OpenStudy (prominer10000):

Last question! :) Now this one really confuses me. (12^-5)^2 -------- = 12 12^-4 Please help! @TheSmartOne

OpenStudy (prominer10000):

@TheSmartOne I see you're helping someone else, don't rush. Just come whenever you can, I can wait. I have all day today :)

OpenStudy (timamits):

please check the question once again.

OpenStudy (thewafflebro):

okay let me see what i can do here

OpenStudy (thewafflebro):

nvm the smart one got it :P

TheSmartOne (thesmartone):

Is the question supposed to be \(\Large\frac{(12^{-5})^2}{12^{-4}} = 12^x\) Is that all equal to 12^x?

OpenStudy (prominer10000):

Yes, the X is the missing exponent.

TheSmartOne (thesmartone):

You need to use this property first: \(\Large (a^b)^c = a^{bc}\) and then you use this property \(\Large \frac{x^y}{x^z} = x^{y-z}\)

OpenStudy (prominer10000):

Ok so, sorry. I'm really confused. So would I add the exponents together or no?

TheSmartOne (thesmartone):

no, you multiply the exponents in the numerator

OpenStudy (prominer10000):

So I do 12 x 5? Or do 12 x 12 x 12 x 12 x 12?

TheSmartOne (thesmartone):

\(\Large (a^b)^c = a^{bc}\) \(\Large (12^{-5})^2 = 12^{??}\)